Select any two integers between -12 and +12 which will become solutions to a system of two equations.
Write two equations that have your two integers as solutions. Show how you built the equations using your integers. Your solution and equations should not be the same as those of other students or the textbook. There are infinite possibilities.
Solve your system of equations by the addition/subtraction method. Make sure you show the necessary 5 steps. Use the example on page 426 of Mathematics in Our World as a guide.

Answer by ankor@dixie-net.com(15746)

Select any two integers between -12 and +12 which will become solutions
to a system of two equations.
:
Choose two integers, x =-5 and y=+10
:
1st equation
3(-5) + 2(10) =
-15 + 20 = 5
3x + 2y = 5
:
2nd equation
6(-5) - 4(10) =
-30 - 40 = -70
6x - 4y = -70
:
Our two equations
3x + 2y = 5
6x - 4y = -70
:
Solve this using the addition method
Multiply the 1st equation by 2, add to the 2nd equation
6x + 4y = 10
6x - 4y = -70
----------------adding eliminates y, find x
12x + 0 = -60
12x = -60
x = -60/12
x = -5
:
Find y using the 1st equation, replace x with -5
3(-5) + 2y = 5
-15 + 2y = 5
2y = 5 + 15
2y = 20
y = 20/2
y = 10
:
You can check the solution again in the 2nd equation